# SysCaps: Language Interfaces for Simulation Surrogates of Complex Systems

### v1

@@ Abstract
Data-driven simulation surrogates help computational scientists study complex systems. They can also help inform impactful policy decisions. We introduce a learning framework for surrogate modeling where language is used to interface with the underlying system being simulated. We call a language description of a system a “system caption”, or SysCap. To address the lack of datasets of paired natural language SysCaps and simulation runs, we use large language models (LLMs) to synthesize high-quality captions. Using our framework, we train multimodal text and timeseries regression models for two real-world simulators of complex energy systems. Our experiments demonstrate the feasibility of designing language interfaces for real-world surrogate models at comparable accuracy to standard baselines. We qualitatively and quantitatively show that SysCaps unlock text-prompt-style surrogate modeling and new generalization abilities beyond what was previously possible. We will release the generated SysCaps datasets and our code to support follow-on studies.
@@ 1 Introduction
Data-driven surrogates enable computational scientists to efficiently predict the results of expensive numerical simulations that run on supercomputers [30, 7]. Surrogates are particularly valuable for emulating simulations of complex energy systems (CES), which model dynamic interactions between humans, earth systems, and infrastructure. Examples of CES include buildings [52, 9, 5], electric vehicle fleets [53], and microgrids [13]. Advancing the science of CES contributes to reducing emissions and accelerating the adoption of clean energy, which is needed to address the impacts of climate change. Surrogate models are not only intended for expert use. Surrogates are also needed to inform highly consequential policy and investment decisions about complex systems made by non-experts in industry and governments [41], such as when planning to build and deploy a new renewable energy system [21].
@@
These types of surrogates perform a fairly standard regression task, predicting simulation output quantities of interest from a) an input system configuration and b) a deployment scenario. For example, we might want to predict a) the amount of energy a particular building will consume given b) a timeseries of weather variables spanning an entire year. In this case, this involves performing long sequence timeseries regression, which classic data-driven regression such as gradient-boosted decision trees have difficulty with [5, 60].
@@
In this work, we aim to explore the design and analysis of language interfaces for such surrogates. Intuitively, a language interface to a surrogate model makes it more accessible, particularly for non-experts, by simplifying how we inspect and alter a complex system’s configuration. Language interfaces are powerful—they ground interactions between humans and machines in the human’s preferred way [51]. The idea of using language to create interfaces for complex data or models is not new [23, 39], and interest has renewed due to the success of large language models (LLMs) and their demonstrated ability to generate high-quality synthetic natural “captions” [46, 12, 36, 22]. Our work defines a “system caption”, or SysCap, as text-based descriptions of knowledge about the system being simulated. In this work, we focus on black-box settings where the only available knowledge about the system is its configuration, found in simulator input files as lists of attributes.
@@
However, it is unclear whether textual inputs, and particularly natural language, is suitable for real-world regression tasks. System attributes are sets of both discrete (categorical, binary, or string) and continuous (numeric) variables, i.e., tabular data. Previous work demonstrated inconclusive evidence when using pretrained language models to do tabular regression from text-encoded inputs, with and without modifications to the architecture [11, 27, 4, 57], motivating further study. Indeed, regression with text-encoded tabular inputs promises multiple advantages. First, it avoids cumbersome feature engineering (e.g., one-hot encodings) for tabular data. Second, it flexibly handles variable-length inputs, such as when certain system attributes are unknown at test time. Third, using text to represent tabular data increases generalizability—intuitively, attribute names and values have semantic information that can be exploited with the help of pretrained language embeddings [22, 57].
@@
Our paper introduces a framework for training lightweight multimodal surrogates for CES with text (for system attributes) and timeseries inputs (for the deployment scenario) and makes contributions towards addressing the following technical challenges:
  • Given the lack of human-labeled natural language descriptions of complex systems, we describe a data collection pipeline that uses an LLM to generate high-quality natural language SysCaps from metadata files of CES. We observe that LLMs possess broad knowledge about CES, thus, minimal prompting is required to produce conversational SysCaps.
@@
  • We introduce a simple and lightweight multimodal surrogate model architecture that a) fuses text embeddings obtained from pretrained language models (LMs) with b) timeseries encoded by a bidirectional sequence encoder to c) regress a timeseries output. We expect this to be insightful for future multimodal text and timeseries studies.
@@
Our experiments are based on two real-world CES simulators of building energy demand and wind farm wake. We rigorously evaluate accuracy and generalization beyond the capabilities of traditional regression approaches by quantifying robustness to variable-length inputs and paraphrasing (e.g., attribute synonyms). We also qualitatively showcase how text interfaces enable flexible handling of missing attributes and rapid design space exploration via the use of captions. As there are no standard benchmarks for comparing surrogate modeling performance for CES, we will open-source all data and code and contribute the generated SysCaps datasets to facilitate future work.
@@ 2 Related Work
Language interfaces for scientific machine learning: A growing body of work is similarly exploring creating language interfaces for advanced scientific machine learning (SciML) tasks, including protein representation learning [56], protein design [32], and activity prediction for drug discovery [47]. LLM-powered natural language interfaces are also being designed for complicated scientific workflows including synchrotron management [38], automated chemistry labs [6], and fluid dynamics workflows [29]. Our work adds to this growing body of literature by studying language interfaces for surrogate models of complex systems.
@@
Large language models for regression: Another line of work asks whether LLMs can perform regression with both language inputs and outputs (numbers encoded as tokens), such as for tabular problems [11] or black-box optimization [48, 33]. These studies use templated text inputs, whereas our work additionally seeks to answer whether natural language is also viable as an interface for regression. Also, our framework directly predicts continuous outputs instead of tokens. Moreover, one study found mixed results compared to simple gradient-boosted tree baselines and difficulty with interpolation [11], raising questions about the effectiveness of this direction. LLMs are also expensive to evaluate, which makes them impractical for scientific applications where the surrogate is intended to be called many times. We only use LLMs to create synthetic training data and instead employ lightweight LMs such as DistilBERT [45] to encode SysCaps.
Text and timeseries multimodal models: Outside of limited prior work on timeseries forecasting with text covariate inputs for taxi demand [44] and financial data [14], most work on multimodal text and timeseries modeling is in audio generation. Here, LLMs have also recently been used to synthesize captions for various tasks [12, 37]. Models for text-guided audio and music generation [1, 26, 31] are pretrained with a contrastive objective that aligns embeddings of captions that describe the audio and music. This differs from our setting, where the SysCaps describes the complex system being simulated, and the timeseries inputs are covariates for the target variable (the simulator outputs).
@@
Knowledge-enhanced PDE surrogates: Numerical simulation of partial differential equations (PDEs) is extremely computationally intensive, and thus a large body of SciML work is focused on developing neural surrogates which are fast to evaluate. Closely related work has recently tried to encode knowledge about the PDE into a neural surrogate to facilitate generalization within and across families of PDEs. Methods that embed equation parameters (i.e., the system attributes) within the architecture to generalize to unseen parameters include CAPE [49] and those explored in Gupta & Brandstetter [19]. Other approaches embed structural knowledge about the PDE equation into the surrogate model architecture [40, 59] or in the loss [42]. Concurrent work has explored “PDE captions” [34, 58], which are a type of SysCaps for neural PDE surrogates where the system knowledge is PDE equations encoded as text.
@@ 3 Problem Statement
Our goal is to learn a surrogate $f:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$ that regresses the outputs of a simulator $F$ directly from its inputs. More formally, we are given a dataset $D$ of pairs of simulator inputs and outputs. The inputs are the system deployment scenario $x_{1:T}\in\mathcal{X}$ (a timeseries) and the tabular system configuration inputs $z\in\mathcal{Z}$. The outputs are a timeseries $y_{1:T}\in\mathcal{Y}$. For simplicity, we consider only univariate timeseries outputs in this work ($y_{t}\in\mathbb{R}$), although the number of timesteps $T$ may be large (potentially thousands of steps), the timeseries inputs $\mathcal{X}$ are multivariate, and the mapping $f$ which approximates the simulator is highly nonlinear.
To summarize, we have a timeseries regression problem modeled as $y_{1:T}=f(x_{1:T},z)$. By conditioning the surrogate on system knowledge $z$, it can potentially generalize to new system configurations. However, learning transferable representations of variable-length, heterogeneous input features such as $z$ is notoriously difficult for deep neural networks, and is a key focus of tabular deep learning (see survey  [3]). In our work, we develop and analyze a framework for learning multimodal surrogates where $z$ is encoded as a text (using templates as well as conversational natural language, Section 4).
@@ 4 Synthesizing System Captions (SysCaps) with LLMs
Figure 1: An example of generating a SysCap by strategically prompting an LLM with system attributes that are represented as key-value pairs. The resulting caption is able to capture the provided attributes in a natural conversational language.
Our work is motivated by the idea that language interfaces for surrogates represent a path towards improving the accessibility of these models for expert and non-expert users, e.g., when using them for downstream system design tasks [51]. In this section, we describe two approaches for converting system attributes into text: key-value templates and natural language.
For the key-value approach, attributes are described as key-value pairs key:value and joined by a separator “||||” (SysCaps-kv). For example, if a simulation has attributes A=1.0 and B=blue, we create the string A:1.0|B:blue. Generating these strings is easy to do and incurs a negligible amount of extra computational overhead. In the natural language approach (SysCaps-nl, Figure 1), attributes are described in a conversational manner, which we believe is more flexible and expressive than key-value captions and thereby more accessible for non-experts. However, we do not have access to large quantities of natural language descriptions for each system and simulation. We avoid the time-consuming task of enlisting domain experts to create this data by instead prompting a powerful LLM to generate synthetic natural language descriptions given attributes. The details of the prompt are provided next. In our work, we use the open-source LLM llama-2-7b-chat [50].
@@
Attribute subset selection: Simulations of real-world systems may have attributes that only weakly correlate with the output quantity of interest, or have a large number of attributes, which can be challenging for deep learning approaches. Since the length of a SysCap is proportional to the number of attributes, the computational burden incurred by text-based encodings of attributes can grow significantly in these cases. In these cases, reducing the number of attributes can be handled with classic feature selection methods such as recursive feature elimination (RFE) [20] or by recommendations from domain experts, as a pre-processing step.
@@ 5 Text and Timeseries Surrogate Model
Figure 2: Building blocks of our surrogate model, $f=h_{\theta}\circ g_{\psi}$, that includes a multimodal encoder, $g_{\psi}$, and a top model, $h_{\theta}$. The multimodal encoder, $g_{\psi}=g_{\psi}^{\mathsf{seq}}\circ g_{\psi}^{\mathsf{text}}$, is a composition of a text encoder, $g_{\psi}^{\mathsf{text}}$, and a bidirectional sequence encoder, $g_{\psi}^{\mathsf{seq}}$, for timeseries inputs.
@@
Multimodal encoder $g_{\psi}$: A text encoder $g_{\psi}^{\mathsf{text}}$ extracts an embedding $\hat{z}$ from a SysCap $z$ with a pretrained LM, then broadcasts and concatenates this text embedding with the timeseries inputs to create a multimodal feature vector for each simulation timestep. These features get processed by a bidirectional sequence encoder $g_{\psi}^{\mathsf{seq}}$ to produce a sequence of time-dependent multimodal features, $e_{1:T}=g_{\psi}^{\mathsf{seq}}(g_{\psi}^{\mathsf{text}}(z),x_{1:T})$ , which are finally used to regress outputs.
Text encoder $g_{\psi}^{\mathsf{text}}$: To encode textual inputs we use pretrained models such as DistilBERT [45] and BERT [10]. We use each model’s default pretrained tokenizer. Tokenized sequences are bracketed by [CLS] and [EOS] tokens, and we use the final activation at the [CLS] token position to produce a text embedding $\hat{z}\in\mathbb{R}^{d}$. Following standard fine-tuning practices, all layers for BERT are fine-tuned while only the last layer of DistilBERT is fine-tuned.
Bidirectional sequence encoder $g_{\psi}^{\mathsf{seq}}$: We broadcast the text embedding $\hat{z}$ to create a sequence of length $T$, $\hat{z}\rightarrow\{\hat{z}_{t}\}_{t=1}^{T}$, and concatenate each $\hat{z}_{t}$ with the timeseries input $x_{1:T}$, $\{\hat{z}_{t};x_{t}\}_{1}^{T}$. This simplifies the task of learning timestep-specific correlations between system attributes $\hat{z}$ and timeseries $x_{1:T}$ in the multimodal encoder $g_{\psi}$. To efficiently embed long timeseries with thousands of timesteps, we explore both bidirectional LSTMs [25] and bidirectional SSMs [16] for $g_{\psi}^{\mathsf{seq}}$. Our bidirectional SSM uses stacks of S4 [18] blocks without downpooling layers. We use the last layer’s hidden states as temporal features $e_{1:T}$ for the top model. If $T=1$ or for non-sequential surrogate models, we instead use an MLP with residual layers (ResNet MLP) to embed each $\{\hat{z}_{t};x_{t}\}$ per-timestep to get $e_{t}$.

### v2

@@ Abstract
Surrogate models are used to predict the behavior of complex energy systems that are too expensive to simulate with traditional numerical methods. Our work introduces the use of language descriptions, which we call “system captions” or SysCaps, to interface with such surrogates. We argue that interacting with surrogates through text, particularly natural language, makes these models more accessible for both experts and non-experts. We introduce a lightweight multimodal text and timeseries regression model and a training pipeline that uses large language models (LLMs) to synthesize high-quality captions from simulation metadata. Our experiments on two real-world simulators of buildings and wind farms show that our SysCaps-augmented surrogates have better accuracy on held-out systems than traditional methods while enjoying new generalization abilities, such as handling semantically related descriptions of the same test system. Additional experiments also highlight the potential of SysCaps to unlock language-driven design space exploration and to regularize training through prompt augmentation.
@@ 1 Introduction
Data-driven surrogates enable computational scientists to efficiently predict the results of expensive numerical simulations that run on supercomputers (Lavin et al., 2021; Carter et al., 2023). Surrogates are particularly valuable for emulating simulations of complex energy systems (CES), which model dynamic interactions between humans, earth systems, and infrastructure. Examples of CES include buildings (Vazquez-Canteli et al., 2019; Dai et al., 2023; Bhavsar et al., 2023), electric vehicle fleets (Vepsäläinen et al., 2019), and microgrids (Du & Li, 2019). Advancing the science of CES contributes to reducing emissions and accelerating the adoption of clean energy, which is needed to address the impacts of climate change.
@@
These models perform a fairly standard regression task, predicting simulation output quantities of interest from a) an input system configuration and b) a deployment scenario. For example, we might want to predict the amount of energy a building will consume given a) a list of building characteristics and b) a weather timeseries spanning an entire year. In this case, this involves performing long sequence timeseries regression, which traditional regression techniques such as gradient-boosted decision trees have difficulty with (Bhavsar et al., 2023; Zhang et al., 2021).
@@
Surrogate models are not only used by experts. Surrogates are also used to inform highly consequential policy and investment decisions about complex systems made by non-experts in industry and governments (Rackauckas & Abdelrehim, 2024), such as when planning to build and deploy a new renewable energy system (Harrison-Atlas et al., 2024). In this work, we design and analyze language interfaces for such surrogates. Intuitively, language interfaces make surrogate models more accessible, particularly for non-experts, by simplifying how we inspect and alter a complex system’s configuration. Language interfaces are powerful—they ground interactions between humans and machines in the human’s preferred way (Vaithilingam et al., 2024). The idea of using language to create interfaces for complex data or models is not new (Hendrix et al., 1978; Quamar et al., 2022), but interest has renewed due to the success of large language models (LLMs) and their demonstrated ability to generate high-quality synthetic natural captions (Schick & Schütze, 2021; Doh et al., 2023; Mei et al., 2023a; Hegselmann et al., 2023). Our work defines a “system caption”, or SysCap, as text-based descriptions of knowledge about the system being simulated. The only available knowledge our work assumes is the system configuration found in simulator metadata files as lists of attributes.
@@
In general, it is unknown whether textual inputs, and particularly natural language inputs, are suitable for real-world tabular regression tasks. Tabular data, such as the system attributes in question, are sets of both discrete (categorical, binary, or string) and continuous (numeric) variables. Previous work demonstrated inconclusive evidence when using pretrained language models to do tabular regression from text-encoded inputs, with and without modifications to the architecture (Dinh et al., 2022; Jablonka et al., 2024; Bellamy et al., 2023; Yan et al., 2024), motivating further study. Regression with text-encoded tabular inputs is promising because a) language is a more intuitive and flexible user interface than traditional encoding strategies (e.g., one-hot encodings), and b) using pretrained language embeddings to encode system attributes unlocks the ability to exploit the semantic information contained in SysCaps to generalize across related systems.
@@
Our paper introduces a framework for training multimodal surrogates for CES with text (for system attributes) and timeseries inputs (for the deployment scenario) and makes contributions towards addressing the following technical challenges:
@@
  • We introduce a simple and lightweight multimodal surrogate model architecture for timeseries regression that a) fuses text embeddings obtained from pretrained language models (LMs) with b) timeseries encoded by a bidirectional sequence encoder. We expect this to be insightful for future multimodal text and timeseries studies.
  • To address the lack of human-labeled natural language descriptions of complex systems, we describe a process that uses LLMs to generate high-quality natural language SysCaps from simulation metadata. Although LLMs have previously been used to generate text captions from metadata (Doh et al., 2023; Mei et al., 2023b), we believe our application to multi-modal surrogate modeling is novel.
@@
Our experiments are based on two real-world CES simulators of buildings and wind farm wake. We rigorously evaluate accuracy on held-out systems and show that SysCaps-augmented surrogates have better accuracy than one-hot baselines. We also show generalization beyond the capabilities of traditional regression approaches enabled by the use of text embeddings, e.g., robustness to replacing attributes names with synonyms in test captions. We qualitatively show that text interfaces unlock system design space exploration via language. As there are no standard benchmarks for comparing surrogate modeling performance for CES, we will open-source all data and code and contribute the generated SysCaps datasets to facilitate future work.
@@ 2 Related Work
Language interfaces for scientific machine learning: An increasing amount of work is exploring language interfaces for advanced scientific machine learning (SciML) models, including protein representation learning (Xu et al., 2023), protein design (Liu et al., 2023b), and activity prediction for drug discovery (Seidl et al., 2023). LLM-powered natural language interfaces are also being designed for complicated scientific workflows including synchrotron management (Potemkin et al., 2023), automated chemistry labs (Bran et al., 2023), and fluid dynamics workflows (Kumar et al., 2023). We add to this body of work by studying language interfaces for lightweight surrogate models.
@@
Large language models for regression: Another line of work asks whether LLMs can perform regression with both text inputs and outputs (numbers encoded as tokens), such as for tabular problems (Dinh et al., 2022) or black-box optimization (Song et al., 2024; Liu et al., 2024). We do not use an LLM to do regression directly, but rather train a lightweight multimodal architecture that predicts continuous outputs instead of tokens. Moreover, one study (Dinh et al., 2022) found mixed results when comparing to simple gradient-boosted tree baselines and highlighted difficulty with interpolation, raising questions about the effectiveness of LLM-based regression.
Multimodal text and timeseries forecasting: Timeseries forecasting aims to predict future values of an input timeseries given past values. Surrogate modeling can be cast as a forecasting problem when the goal is to train a model to emulate a dynamical system and predict its future behavior (e.g., predicting future energy demand from past energy usage). Previous work has explored multimodal timeseries forecasting where auxiliary text data is introduced as covariates to improve the forecasting accuracy (Rodrigues et al., 2019; Emami et al., 2023a; Jin et al., 2024). Notably, Time-LLM (Jin et al., 2024) “reprograms” an LLM to process both text prompts and timeseries. However, our timeseries regression setting (Section 3) differs in that our surrogates are trained to map simulator inputs (e.g., building characteristics and a weather timeseries) to simulator outputs (e.g., energy usage). In our problem, models critically depend on the system information encoded as text, whereas in Time-LLM the text only contains auxiliary information that slightly improves forecasting accuracy. This critical dependence partially motivates our development of a custom architecture for fusing text and timeseries embeddings. Recent evidence (Merrill et al., 2024; Tan et al., 2024) also brings into question whether LLMs are useful at all for reasoning about temporal data; thus, research on lightweight multimodal models is warranted.
Multimodal text and timeseries contrastive pretraining: Various efforts have explored contrastive pretraining objectives for modeling text and timeseries data (Agostinelli et al., 2023; Huang et al., 2022; Liu et al., 2023a; Zhou et al., 2023). For example, Agostinelli et al. (2023) aligns embeddings of captions that describe the audio. In our setting, captions describe system attributes, and the timeseries inputs are exogenous simulator inputs. Here, the text and timeseries inputs share no information with which to learn a shared embedding space.
@@
Knowledge-enhanced PDE surrogates: Numerical simulation of partial differential equations (PDEs) is extremely computationally intensive, and thus a large body of SciML work is focused on developing neural PDE surrogates which are efficient to evaluate. Recent work has tried to encode knowledge about the physical system under study (the PDE) into a surrogate to facilitate generalization within and across families of PDEs. Specifically, methods that embed equation parameters (i.e., the system attributes) within the architecture to generalize to unseen parameters include CAPE (Takamoto et al., 2023) and those explored in Gupta & Brandstetter (2022). Others embed structural knowledge about the PDE into the surrogate model architecture (Rackauckas et al., 2020; Ye et al., 2024) or the loss (Raissi et al., 2019). Concurrent work has explored “PDE captions” (Lorsung et al., 2024; Yang et al., 2024), which are a type of SysCaps for neural PDE surrogates where the system knowledge is PDE equations encoded as text.
@@ 3 Problem Statement
Our goal is to learn a surrogate $f:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$ that regresses the outputs of a simulator $F$ directly from its inputs. We are given a dataset $D$ of pairs of simulator inputs and outputs. The inputs are a deployment scenario $x_{1:T}\in\mathcal{X}$ (a timeseries) and the tabular system attributes $z\in\mathcal{Z}$. The outputs are a timeseries $y_{1:T}\in\mathcal{Y}$. For simplicity, we consider only univariate timeseries outputs in this work ($y_{t}\in\mathbb{R}$). However, the number of timesteps $T$ may be large ( thousands of steps), the timeseries inputs $\mathcal{X}$ are multivariate, and the mapping $f$ which approximates the simulator is highly nonlinear.
To summarize, we have a timeseries regression problem modeled as $y_{1:T}=f(x_{1:T},z)$. By conditioning the surrogate on system knowledge $z$, it can potentially generalize to new system configurations. However, learning transferable representations of variable-length, heterogeneous input features such as $z$ is notoriously difficult for deep neural networks, and is a key focus of tabular deep learning (see survey  (Badaro et al., 2023)). In our work, we develop and analyze a framework for learning multimodal surrogates where $z$ is encoded as text.
@@ 4 Synthesizing System Captions (SysCaps) with LLMs
Figure 1: Our pipeline for augmenting multimodal simulation surrogates with language interfaces using “system captions”, or SysCaps. SysCaps are text descriptions of knowledge about the system being simulated. In our work, the SysCaps describe the system’s characteristics, as found in simulation metadata files. During training (a), we create paired datasets of temporal simulator inputs with key-value template SysCaps or LLM-generated natural language SysCaps. At test time (b), we prompt the surrogate model with one or more key-value template captions or natural language captions. LLMs are only used to generate synthetic training data; we use a lightweight BERT-style text encoder and an efficient long-sequence encoder to keep the computational cost of our surrogate low.
Our work is motivated by the idea that language interfaces for surrogates represent a path towards improving the accessibility of these models for expert and non-expert users, e.g., when using them for downstream system design tasks (Vaithilingam et al., 2024). Our proposed framework for augmenting surrogates with language interfaces is visualized in Figure 1. During training, we create SysCaps out of system attributes specified in simulation metadata. To create large amounts of synthetic natural language SysCaps, we use LLMs. The ultimate goal is to enable scientists to “chat" with the multi-modal surrogate model at test time via text prompting. In this section, we describe two approaches for converting system attributes into text: key-value templates and natural language.
For the key-value approach, attributes are described as key-value pairs key:value and joined by a separator “||||” (SysCaps-kv). For example, if a simulation has attributes A=1.0 and B=blue, we create the string A:1.0|B:blue. Generating these strings is easy to do and incurs a negligible amount of extra computational overhead. In the natural language approach (SysCaps-nl, Figure 1), attributes are described in a conversational manner, which we believe is more flexible and expressive than key-value captions and thereby more accessible for non-experts. However, we do not have access to large quantities of natural language descriptions for each system and simulation. We avoid the time-consuming task of enlisting domain experts to create this data by instead prompting a powerful LLM to generate synthetic natural language descriptions given attributes. In our work, we use the open-source LLM llama-2-7b-chat (Touvron et al., 2023). The details of the prompt are provided next.
@@
Attribute subset selection: Simulations of real-world systems may have attributes that only weakly correlate with the output quantity of interest, or have a large number of attributes, which can be challenging for deep learning approaches. Since the length of a SysCap is proportional to the number of attributes, the computational burden incurred by text-based encodings of attributes can grow significantly in these cases. In these cases, reducing the number of attributes can be handled with classic feature selection methods such as recursive feature elimination (RFE) (Guyon et al., 2002) or by recommendations from domain experts, as a pre-processing step.
@@ 5 Text and Timeseries Surrogate Model
Figure 2: Building blocks of our surrogate model, $f=h_{\theta}\circ g_{\psi}$, that includes a multimodal encoder, $g_{\psi}$, and a top model, $h_{\theta}$. The multimodal encoder, $g_{\psi}=g_{\psi}^{\mathsf{seq}}\circ g_{\psi}^{\mathsf{text}}$, is a composition of a text encoder, $g_{\psi}^{\mathsf{text}}$, and a bidirectional sequence encoder, $g_{\psi}^{\mathsf{seq}}$, for timeseries inputs. The text embedding vector $\hat{z}$ is broadcasted (dashed lines) to create a sequence that is concatenated with the timeseries input. This multimodal sequence is the input to the sequence encoder.
@@
Multimodal encoder $g_{\psi}$: A text encoder $g_{\psi}^{\mathsf{text}}$ extracts an embedding $\hat{z}$ from a SysCap $z$ then broadcasts and concatenates this embedding with the timeseries inputs to create a sequence of multimodal feature vectors. These features get processed by a bidirectional sequence encoder $g_{\psi}^{\mathsf{seq}}$ to produce a sequence of time-dependent fused multimodal features $e_{1:T}$, $e_{1:T}=g_{\psi}^{\mathsf{seq}}(g_{\psi}^{\mathsf{text}}(z),x_{1:T})$ , which are finally used to regress outputs.
Text encoder $g_{\psi}^{\mathsf{text}}$: To encode textual inputs we use pretrained BERT (Devlin et al., 2018) and DistilBERT (Sanh et al., 2019) models that are relatively more efficient than LLMs. We use the model’s default pretrained tokenizer. Tokenized sequences are bracketed by [CLS] and [EOS] tokens, and we use the final activation at the [CLS] token position to produce a text embedding $\hat{z}\in\mathbb{R}^{d}$. Following standard fine-tuning practices, all layers for BERT are fine-tuned while only the last layer of DistilBERT is fine-tuned.
Bidirectional sequence encoder $g_{\psi}^{\mathsf{seq}}$: We broadcast the text embedding $\hat{z}$ to create a sequence of length $T$, $\hat{z}\rightarrow\{\hat{z}_{t}\}_{t=1}^{T}$, and concatenate each $\hat{z}_{t}$ with the timeseries input $x_{1:T}$, $\{\hat{z}_{t};x_{t}\}_{1}^{T}$. This simplifies the task of learning timestep-specific correlations between system attributes $\hat{z}$ and timeseries $x_{1:T}$ in the multimodal encoder $g_{\psi}$. To efficiently embed long timeseries with thousands of timesteps, we explore both bidirectional LSTMs (Hochreiter & Schmidhuber, 1997) and bidirectional SSMs (Goel et al., 2022) for $g_{\psi}^{\mathsf{seq}}$. Our bidirectional SSM uses stacks of S4 blocks (Gu et al., 2021) without downpooling layers. We use the last layer’s hidden states as temporal features $e_{1:T}$ for the top model. If $T=1$ or for non-sequential surrogate models, we instead use an MLP with residual layers (ResNet MLP) to embed each $\{\hat{z}_{t};x_{t}\}$ per-timestep to get $e_{t}$.